# SKADA-Bench: Benchmarking Unsupervised Domain Adaptation Methods with Realistic Validation

### v1

@@ D.1 Results per datasets
Train Src 0.34 $\pm$ 0.02 0.63 $\pm$ 0.01 0.7 $\pm$ 0.01 0.46 $\pm$ 0.02 0.6 $\pm$ 0.01 0.63 $\pm$ 0.01 0.5 $\pm$ 0.02 0.36 $\pm$ 0.02 0.71 $\pm$ 0.01 0.64 $\pm$ 0.01 0.38 $\pm$ 0.01 0.74 $\pm$ 0.02 0.56 $\pm$ 0.02 10.08
Train Tgt 0.71 $\pm$ 0.02 0.92 $\pm$ 0.01 0.84 $\pm$ 0.01 0.73 $\pm$ 0.03 0.91 $\pm$ 0.01 0.83 $\pm$ 0.0 0.73 $\pm$ 0.01 0.7 $\pm$ 0.01 0.83 $\pm$ 0.01 0.75 $\pm$ 0.02 0.7 $\pm$ 0.01 0.91 $\pm$ 0.01 0.8 $\pm$ 0.01 1.00
Reweighting Dens. RW 0.34 $\pm$ 0.02 0.63 $\pm$ 0.02 0.7 $\pm$ 0.01 0.46 $\pm$ 0.02 0.56 $\pm$ 0.07 0.63 $\pm$ 0.01 0.5 $\pm$ 0.02 0.36 $\pm$ 0.02 0.71 $\pm$ 0.01 0.64 $\pm$ 0.01 0.36 $\pm$ 0.04 0.74 $\pm$ 0.02 0.55 $\pm$ 0.02 9.62
Disc. RW 0.32 $\pm$ 0.02 0.56 $\pm$ 0.03 0.7 $\pm$ 0.01 0.43 $\pm$ 0.01 0.53 $\pm$ 0.01 0.56 $\pm$ 0.02 0.52 $\pm$ 0.02 0.35 $\pm$ 0.01 0.72 $\pm$ 0.01 0.63 $\pm$ 0.01 0.36 $\pm$ 0.01 0.72 $\pm$ 0.02 0.53 $\pm$ 0.01 14.33
Gauss. RW 0.22 $\pm$ 0.01 0.43 $\pm$ 0.01 0.6 $\pm$ 0.02 0.36 $\pm$ 0.02 0.46 $\pm$ 0.02 0.49 $\pm$ 0.02 0.41 $\pm$ 0.02 0.25 $\pm$ 0.02 0.61 $\pm$ 0.01 0.53 $\pm$ 0.01 0.26 $\pm$ 0.01 0.67 $\pm$ 0.01 0.44 $\pm$ 0.02 18.83
KLIEP 0.34 $\pm$ 0.02 0.63 $\pm$ 0.02 0.7 $\pm$ 0.01 0.46 $\pm$ 0.02 0.6 $\pm$ 0.01 0.63 $\pm$ 0.02 0.51 $\pm$ 0.02 0.35 $\pm$ 0.02 0.71 $\pm$ 0.01 0.63 $\pm$ 0.01 0.38 $\pm$ 0.01 0.74 $\pm$ 0.03 0.56 $\pm$ 0.02 10.08
KMM 0.31 $\pm$ 0.03 0.62 $\pm$ 0.02 0.69 $\pm$ 0.01 0.48 $\pm$ 0.03 0.6 $\pm$ 0.01 0.65 $\pm$ 0.02 0.48 $\pm$ 0.01 0.35 $\pm$ 0.02 0.71 $\pm$ 0.01 0.59 $\pm$ 0.01 0.36 $\pm$ 0.01 0.73 $\pm$ 0.02 0.55 $\pm$ 0.02 14.00
NN RW 0.34 $\pm$ 0.02 0.63 $\pm$ 0.01 0.7 $\pm$ 0.01 0.46 $\pm$ 0.02 0.59 $\pm$ 0.01 0.62 $\pm$ 0.02 0.51 $\pm$ 0.02 0.35 $\pm$ 0.02 0.71 $\pm$ 0.01 0.63 $\pm$ 0.01 0.37 $\pm$ 0.01 0.74 $\pm$ 0.02 0.55 $\pm$ 0.02 12.50
MMDTarS 0.34 $\pm$ 0.02 0.63 $\pm$ 0.02 0.63 $\pm$ 0.17 0.46 $\pm$ 0.01 0.6 $\pm$ 0.01 0.63 $\pm$ 0.01 0.51 $\pm$ 0.02 0.36 $\pm$ 0.02 0.71 $\pm$ 0.01 0.64 $\pm$ 0.01 0.38 $\pm$ 0.01 0.74 $\pm$ 0.02 0.55 $\pm$ 0.03 8.67
Mapping CORAL 0.39 $\pm$ 0.02 0.64 $\pm$ 0.01 0.7 $\pm$ 0.01 0.48 $\pm$ 0.02 0.59 $\pm$ 0.01 0.6 $\pm$ 0.01 0.54 $\pm$ 0.02 0.38 $\pm$ 0.01 0.72 $\pm$ 0.01 0.64 $\pm$ 0.01 0.43 $\pm$ 0.01 0.75 $\pm$ 0.02 0.57 $\pm$ 0.01 6.62
MapOT 0.35 $\pm$ 0.02 0.54 $\pm$ 0.02 0.64 $\pm$ 0.01 0.41 $\pm$ 0.02 0.51 $\pm$ 0.01 0.54 $\pm$ 0.01 0.45 $\pm$ 0.02 0.37 $\pm$ 0.02 0.64 $\pm$ 0.02 0.6 $\pm$ 0.02 0.41 $\pm$ 0.01 0.67 $\pm$ 0.02 0.51 $\pm$ 0.02 12.67
EntOT 0.38 $\pm$ 0.02 0.65 $\pm$ 0.01 0.71 $\pm$ 0.01 0.51 $\pm$ 0.01 0.62 $\pm$ 0.02 0.66 $\pm$ 0.02 0.55 $\pm$ 0.02 0.38 $\pm$ 0.01 0.72 $\pm$ 0.02 0.64 $\pm$ 0.01 0.42 $\pm$ 0.01 0.73 $\pm$ 0.01 0.58 $\pm$ 0.01 4.92
ClassRegOT 0.33 $\pm$ 0.04 0.61 $\pm$ 0.07 0.67 $\pm$ 0.05 0.44 $\pm$ 0.05 0.54 $\pm$ 0.05 0.62 $\pm$ 0.05 0.51 $\pm$ 0.06 0.34 $\pm$ 0.03 0.66 $\pm$ 0.02 0.62 $\pm$ 0.03 0.38 $\pm$ 0.04 0.69 $\pm$ 0.05 0.53 $\pm$ 0.04 13.67
LinOT 0.39 $\pm$ 0.02 0.64 $\pm$ 0.02 0.7 $\pm$ 0.01 0.5 $\pm$ 0.02 0.59 $\pm$ 0.01 0.61 $\pm$ 0.01 0.54 $\pm$ 0.02 0.38 $\pm$ 0.01 0.72 $\pm$ 0.01 0.64 $\pm$ 0.01 0.43 $\pm$ 0.01 0.75 $\pm$ 0.03 0.57 $\pm$ 0.01 5.67
MMD-LS 0.34 $\pm$ 0.02 0.6 $\pm$ 0.04 0.68 $\pm$ 0.03 0.46 $\pm$ 0.01 0.56 $\pm$ 0.09 0.58 $\pm$ 0.1 0.5 $\pm$ 0.02 0.32 $\pm$ 0.06 0.71 $\pm$ 0.01 0.63 $\pm$ 0.01 0.36 $\pm$ 0.05 0.73 $\pm$ 0.03 0.54 $\pm$ 0.04 12.79
Subspace JPCA 0.33 $\pm$ 0.02 0.63 $\pm$ 0.01 0.7 $\pm$ 0.01 0.44 $\pm$ 0.02 0.58 $\pm$ 0.01 0.61 $\pm$ 0.02 0.18 $\pm$ 0.02 0.17 $\pm$ 0.01 0.65 $\pm$ 0.01 0.48 $\pm$ 0.05 0.19 $\pm$ 0.05 0.74 $\pm$ 0.02 0.47 $\pm$ 0.02 13.54
SA 0.39 $\pm$ 0.02 0.64 $\pm$ 0.01 0.71 $\pm$ 0.01 0.47 $\pm$ 0.02 0.6 $\pm$ 0.01 0.63 $\pm$ 0.01 0.52 $\pm$ 0.01 0.38 $\pm$ 0.01 0.72 $\pm$ 0.01 0.64 $\pm$ 0.01 0.39 $\pm$ 0.02 0.75 $\pm$ 0.03 0.57 $\pm$ 0.01 6.38
TCA 0.02 $\pm$ 0.0 0.01 $\pm$ 0.0 0.02 $\pm$ 0.0 0.02 $\pm$ 0.01 0.02 $\pm$ 0.0 0.02 $\pm$ 0.01 0.01 $\pm$ 0.01 NA NA 0.04 $\pm$ 0.01 NA NA 0.02 $\pm$ 0.0 21.00
TSL 0.09 $\pm$ 0.02 0.17 $\pm$ 0.05 0.32 $\pm$ 0.01 0.11 $\pm$ 0.01 0.15 $\pm$ 0.05 0.16 $\pm$ 0.04 0.16 $\pm$ 0.03 0.12 $\pm$ 0.01 0.35 $\pm$ 0.01 0.26 $\pm$ 0.09 0.13 $\pm$ 0.01 0.37 $\pm$ 0.05 0.2 $\pm$ 0.03 20.00
Other JDOT 0.35 $\pm$ 0.02 0.5 $\pm$ 0.01 0.62 $\pm$ 0.02 0.42 $\pm$ 0.03 0.54 $\pm$ 0.01 0.55 $\pm$ 0.02 0.48 $\pm$ 0.01 0.37 $\pm$ 0.02 0.64 $\pm$ 0.02 0.59 $\pm$ 0.02 0.4 $\pm$ 0.01 0.66 $\pm$ 0.01 0.51 $\pm$ 0.02 9.08
OTLabelProp 0.32 $\pm$ 0.03 0.57 $\pm$ 0.01 0.66 $\pm$ 0.02 0.5 $\pm$ 0.02 0.63 $\pm$ 0.01 0.63 $\pm$ 0.02 0.53 $\pm$ 0.01 0.39 $\pm$ 0.01 0.7 $\pm$ 0.02 0.63 $\pm$ 0.02 0.43 $\pm$ 0.01 0.73 $\pm$ 0.02 0.56 $\pm$ 0.02 9.67
@@
Table D.4: Accuracy score for 20NewsGroups dataset for each shift compared for all the methods. A white color means the method does not increase the performance compared to Train Src (Train on the source). Green indicates that the performance improved with the DA methods. The darker the color, the more significant the change.
@@
Train Src 0.52 $\pm$ 0.03 0.56 $\pm$ 0.05 0.54 $\pm$ 0.05 0.7 $\pm$ 0.01 0.52 $\pm$ 0.05 0.71 $\pm$ 0.03 0.59 $\pm$ 0.04 18.33
@@
NN RW 0.52 $\pm$ 0.06 0.57 $\pm$ 0.06 0.54 $\pm$ 0.03 0.7 $\pm$ 0.01 0.52 $\pm$ 0.06 0.71 $\pm$ 0.03 0.59 $\pm$ 0.04 17.08
MMDTarS 0.52 $\pm$ 0.03 0.57 $\pm$ 0.05 0.54 $\pm$ 0.05 0.7 $\pm$ 0.01 0.52 $\pm$ 0.06 0.71 $\pm$ 0.03 0.59 $\pm$ 0.04 14.58
Mapping CORAL 0.61 $\pm$ 0.03 0.79 $\pm$ 0.02 0.62 $\pm$ 0.01 0.76 $\pm$ 0.01 0.78 $\pm$ 0.01 0.78 $\pm$ 0.02 0.73 $\pm$ 0.02 11.83
MapOT 0.67 $\pm$ 0.02 0.83 $\pm$ 0.01 0.68 $\pm$ 0.02 0.78 $\pm$ 0.02 0.83 $\pm$ 0.01 0.79 $\pm$ 0.02 0.76 $\pm$ 0.02 7.08
@@
ClassRegOT 0.97 $\pm$ 0.02 0.99 $\pm$ 0.01 0.98 $\pm$ 0.0 0.94 $\pm$ 0.04 0.99 $\pm$ 0.01 0.92 $\pm$ 0.06 0.96 $\pm$ 0.02 2.50
LinOT 0.72 $\pm$ 0.02 0.92 $\pm$ 0.0 0.72 $\pm$ 0.01 0.79 $\pm$ 0.11 0.91 $\pm$ 0.01 0.83 $\pm$ 0.07 0.82 $\pm$ 0.04 6.83
MMD-LS 0.99 $\pm$ 0.0 0.98 $\pm$ 0.0 0.99 $\pm$ 0.0 0.94 $\pm$ 0.01 0.98 $\pm$ 0.0 0.93 $\pm$ 0.01 0.97 $\pm$ 0.01 2.50
Subspace JPCA 0.69 $\pm$ 0.03 0.85 $\pm$ 0.06 0.71 $\pm$ 0.03 0.75 $\pm$ 0.03 0.87 $\pm$ 0.05 0.76 $\pm$ 0.02 0.77 $\pm$ 0.04 6.33
SA 0.9 $\pm$ 0.08 0.91 $\pm$ 0.08 0.87 $\pm$ 0.06 0.87 $\pm$ 0.01 0.85 $\pm$ 0.06 0.87 $\pm$ 0.01 0.88 $\pm$ 0.05 4.83
TCA 0.58 $\pm$ 0.13 0.58 $\pm$ 0.05 0.56 $\pm$ 0.1 0.6 $\pm$ 0.07 0.56 $\pm$ 0.02 0.56 $\pm$ 0.01 0.57 $\pm$ 0.06 17.25
TSL 0.75 $\pm$ 0.04 0.64 $\pm$ 0.05 0.74 $\pm$ 0.04 0.68 $\pm$ 0.02 0.62 $\pm$ 0.05 0.68 $\pm$ 0.03 0.68 $\pm$ 0.04 11.83
Other JDOT 0.67 $\pm$ 0.02 0.84 $\pm$ 0.01 0.68 $\pm$ 0.02 0.79 $\pm$ 0.02 0.84 $\pm$ 0.01 0.8 $\pm$ 0.01 0.77 $\pm$ 0.01 6.33
@@
DASVM 0.57 $\pm$ 0.11 0.75 $\pm$ 0.02 0.52 $\pm$ 0.06 0.72 $\pm$ 0.01 0.81 $\pm$ 0.03 0.73 $\pm$ 0.03 0.68 $\pm$ 0.04 14.00
Table D.5: Accuracy score for AmazonReview dataset for each shift compared for all the methods. A white color means the method does not increase the performance compared to Train Src (Train on the source). Green indicates that the performance improved with the DA methods. The darker the color, the more significant the change.

### v2

@@ E.1 Results per datasets
Train Src 0.55 $\pm$ 0.04 0.51 $\pm$ 0.03 0.5 $\pm$ 0.02 0.9 $\pm$ 0.02 0.49 $\pm$ 0.01 0.59 $\pm$ 0.02 8.40
Train Tgt 0.94 $\pm$ 0.03 0.95 $\pm$ 0.01 0.78 $\pm$ 0.02 0.95 $\pm$ 0.02 0.77 $\pm$ 0.01 0.88 $\pm$ 0.02 1.00
Reweighting Dens. RW 0.49 $\pm$ 0.05 0.51 $\pm$ 0.04 0.5 $\pm$ 0.02 0.85 $\pm$ 0.09 0.49 $\pm$ 0.01 0.57 $\pm$ 0.04 11.40
Disc. RW 0.56 $\pm$ 0.06 0.49 $\pm$ 0.03 0.49 $\pm$ 0.02 0.89 $\pm$ 0.03 0.46 $\pm$ 0.01 0.58 $\pm$ 0.03 13.30
Gauss. RW 0.34 $\pm$ 0.03 0.11 $\pm$ 0.04 0.05 $\pm$ 0.02 0.38 $\pm$ 0.04 0.11 $\pm$ 0.03 0.2 $\pm$ 0.03 19.80
KLIEP 0.55 $\pm$ 0.05 0.5 $\pm$ 0.04 0.49 $\pm$ 0.02 0.91 $\pm$ 0.02 0.48 $\pm$ 0.01 0.59 $\pm$ 0.03 10.60
KMM 0.56 $\pm$ 0.06 0.54 $\pm$ 0.04 0.49 $\pm$ 0.03 0.87 $\pm$ 0.03 0.47 $\pm$ 0.02 0.59 $\pm$ 0.03 10.20
NN RW 0.54 $\pm$ 0.04 0.5 $\pm$ 0.03 0.49 $\pm$ 0.02 0.91 $\pm$ 0.01 0.48 $\pm$ 0.01 0.58 $\pm$ 0.02 12.30
MMDTarS 0.53 $\pm$ 0.04 0.52 $\pm$ 0.04 0.5 $\pm$ 0.03 0.81 $\pm$ 0.14 0.49 $\pm$ 0.01 0.57 $\pm$ 0.05 10.00
Mapping CORAL 0.53 $\pm$ 0.04 0.53 $\pm$ 0.03 0.5 $\pm$ 0.03 0.91 $\pm$ 0.03 0.5 $\pm$ 0.01 0.6 $\pm$ 0.03 6.20
MapOT 0.47 $\pm$ 0.03 0.5 $\pm$ 0.04 0.47 $\pm$ 0.01 0.85 $\pm$ 0.03 0.43 $\pm$ 0.01 0.54 $\pm$ 0.03 13.30
EntOT 0.52 $\pm$ 0.03 0.56 $\pm$ 0.02 0.5 $\pm$ 0.03 0.88 $\pm$ 0.02 0.46 $\pm$ 0.01 0.58 $\pm$ 0.02 11.80
ClassRegOT 0.58 $\pm$ 0.07 0.59 $\pm$ 0.04 0.51 $\pm$ 0.02 0.84 $\pm$ 0.04 0.25 $\pm$ 0.25 0.55 $\pm$ 0.08 9.20
LinOT 0.54 $\pm$ 0.05 0.56 $\pm$ 0.05 0.49 $\pm$ 0.02 0.9 $\pm$ 0.02 0.49 $\pm$ 0.01 0.6 $\pm$ 0.03 8.90
MMD-LS 0.55 $\pm$ 0.04 0.51 $\pm$ 0.04 0.49 $\pm$ 0.03 0.9 $\pm$ 0.02 0.48 $\pm$ 0.01 0.59 $\pm$ 0.03 8.00
Subspace JPCA 0.53 $\pm$ 0.04 0.51 $\pm$ 0.02 0.49 $\pm$ 0.02 0.89 $\pm$ 0.01 0.48 $\pm$ 0.01 0.58 $\pm$ 0.02 14.10
SA 0.42 $\pm$ 0.17 0.5 $\pm$ 0.03 0.5 $\pm$ 0.02 0.84 $\pm$ 0.09 0.49 $\pm$ 0.02 0.55 $\pm$ 0.07 12.10
TCA 0.03 $\pm$ 0.01 0.03 $\pm$ 0.0 0.04 $\pm$ 0.0 0.04 $\pm$ 0.0 0.04 $\pm$ 0.0 0.03 $\pm$ 0.0 20.80
TSL 0.55 $\pm$ 0.04 0.51 $\pm$ 0.04 0.49 $\pm$ 0.03 0.9 $\pm$ 0.02 0.49 $\pm$ 0.01 0.59 $\pm$ 0.03 9.20
Other JDOT 0.64 $\pm$ 0.05 0.63 $\pm$ 0.02 0.5 $\pm$ 0.02 0.77 $\pm$ 0.04 0.49 $\pm$ 0.01 0.61 $\pm$ 0.03 6.20
OTLabelProp 0.58 $\pm$ 0.03 0.63 $\pm$ 0.02 0.49 $\pm$ 0.04 0.89 $\pm$ 0.02 NA 0.65 $\pm$ 0.03 8.75
@@
Table 8: Accuracy score for Office31 dataset for each shift compared for all the deep DA methods. A white color means the method does not increase the performance compared to Train Src (Train on the source). Green indicates that the performance improved with the DA methods. The darker the color, the more significant the change.
@@
Train Src 0.72 0.75 0.61 0.94 0.63 0.99 0.77 4.17
Train Tgt 0.99 1.0 0.87 0.99 0.88 0.99 0.95 1.25
DANN 0.75 0.8 0.64 0.96 0.62 0.98 0.79 3.75
DeepCORAL 0.77 0.77 0.61 0.98 0.63 0.99 0.79 3.33
DeepJDOT 0.79 0.8 0.68 0.97 0.69 1.0 0.82 2.17
Table 9: Accuracy score for OfficeHome dataset for each shift compared for all the methods. A white color means the method does not increase the performance compared to Train Src (Train on the source). Green indicates that the performance improved with the DA methods. The darker the color, the more significant the change.
[…]
Train Src 0.34 $\pm$ 0.02 0.63 $\pm$ 0.01 0.7 $\pm$ 0.01 0.46 $\pm$ 0.02 0.6 $\pm$ 0.01 0.63 $\pm$ 0.01 0.5 $\pm$ 0.02 0.36 $\pm$ 0.02 0.71 $\pm$ 0.01 0.63 $\pm$ 0.01 0.38 $\pm$ 0.01 0.74 $\pm$ 0.02 0.56 $\pm$ 0.01 10.42
Train Tgt 0.71 $\pm$ 0.02 0.92 $\pm$ 0.01 0.84 $\pm$ 0.01 0.73 $\pm$ 0.03 0.91 $\pm$ 0.01 0.83 $\pm$ 0.01 0.72 $\pm$ 0.01 0.7 $\pm$ 0.01 0.83 $\pm$ 0.01 0.75 $\pm$ 0.02 0.7 $\pm$ 0.01 0.91 $\pm$ 0.01 0.8 $\pm$ 0.01 1.00
Reweighting Dens. RW 0.34 $\pm$ 0.02 0.63 $\pm$ 0.01 0.7 $\pm$ 0.01 0.47 $\pm$ 0.02 0.58 $\pm$ 0.06 0.63 $\pm$ 0.01 0.5 $\pm$ 0.02 0.36 $\pm$ 0.02 0.71 $\pm$ 0.01 0.64 $\pm$ 0.01 0.37 $\pm$ 0.03 0.73 $\pm$ 0.05 0.55 $\pm$ 0.02 9.71
Disc. RW 0.32 $\pm$ 0.01 0.56 $\pm$ 0.03 0.71 $\pm$ 0.02 0.42 $\pm$ 0.02 0.52 $\pm$ 0.01 0.57 $\pm$ 0.02 0.53 $\pm$ 0.01 0.35 $\pm$ 0.02 0.72 $\pm$ 0.01 0.63 $\pm$ 0.01 0.37 $\pm$ 0.01 0.73 $\pm$ 0.01 0.54 $\pm$ 0.02 13.00
Gauss. RW 0.22 $\pm$ 0.01 0.43 $\pm$ 0.01 0.59 $\pm$ 0.02 0.35 $\pm$ 0.01 0.46 $\pm$ 0.03 0.48 $\pm$ 0.02 0.41 $\pm$ 0.02 0.25 $\pm$ 0.02 0.61 $\pm$ 0.01 0.53 $\pm$ 0.01 0.26 $\pm$ 0.02 0.67 $\pm$ 0.02 0.44 $\pm$ 0.02 18.08
KLIEP 0.34 $\pm$ 0.02 0.63 $\pm$ 0.01 0.7 $\pm$ 0.01 0.46 $\pm$ 0.02 0.6 $\pm$ 0.01 0.63 $\pm$ 0.01 0.5 $\pm$ 0.02 0.36 $\pm$ 0.02 0.71 $\pm$ 0.01 0.64 $\pm$ 0.01 0.38 $\pm$ 0.01 0.74 $\pm$ 0.03 0.56 $\pm$ 0.02 9.00
KMM 0.32 $\pm$ 0.02 0.62 $\pm$ 0.01 0.7 $\pm$ 0.01 0.45 $\pm$ 0.07 0.61 $\pm$ 0.02 0.64 $\pm$ 0.02 0.48 $\pm$ 0.01 0.34 $\pm$ 0.02 0.7 $\pm$ 0.01 0.58 $\pm$ 0.01 0.35 $\pm$ 0.04 0.73 $\pm$ 0.01 0.54 $\pm$ 0.02 13.96
NN RW 0.34 $\pm$ 0.02 0.63 $\pm$ 0.01 0.7 $\pm$ 0.02 0.45 $\pm$ 0.02 0.59 $\pm$ 0.01 0.61 $\pm$ 0.01 0.5 $\pm$ 0.02 0.35 $\pm$ 0.02 0.71 $\pm$ 0.01 0.63 $\pm$ 0.01 0.37 $\pm$ 0.01 0.74 $\pm$ 0.02 0.55 $\pm$ 0.02 12.96
MMDTarS 0.34 $\pm$ 0.02 0.63 $\pm$ 0.01 0.7 $\pm$ 0.01 0.46 $\pm$ 0.02 0.61 $\pm$ 0.01 0.63 $\pm$ 0.01 0.5 $\pm$ 0.02 0.36 $\pm$ 0.02 0.71 $\pm$ 0.01 0.64 $\pm$ 0.01 0.38 $\pm$ 0.01 0.74 $\pm$ 0.03 0.56 $\pm$ 0.02 7.71
Mapping CORAL 0.39 $\pm$ 0.02 0.64 $\pm$ 0.01 0.7 $\pm$ 0.0 0.48 $\pm$ 0.02 0.58 $\pm$ 0.0 0.61 $\pm$ 0.01 0.54 $\pm$ 0.02 0.38 $\pm$ 0.01 0.72 $\pm$ 0.01 0.64 $\pm$ 0.01 0.43 $\pm$ 0.01 0.75 $\pm$ 0.02 0.57 $\pm$ 0.01 6.62
MapOT 0.36 $\pm$ 0.03 0.53 $\pm$ 0.01 0.64 $\pm$ 0.01 0.41 $\pm$ 0.02 0.52 $\pm$ 0.01 0.53 $\pm$ 0.01 0.45 $\pm$ 0.01 0.37 $\pm$ 0.02 0.64 $\pm$ 0.01 0.59 $\pm$ 0.02 0.41 $\pm$ 0.01 0.67 $\pm$ 0.02 0.51 $\pm$ 0.02 12.42
EntOT 0.38 $\pm$ 0.02 0.65 $\pm$ 0.01 0.72 $\pm$ 0.01 0.5 $\pm$ 0.01 0.61 $\pm$ 0.02 0.66 $\pm$ 0.02 0.55 $\pm$ 0.02 0.38 $\pm$ 0.01 0.71 $\pm$ 0.02 0.65 $\pm$ 0.01 0.42 $\pm$ 0.01 0.73 $\pm$ 0.01 0.58 $\pm$ 0.01 5.00
ClassRegOT 0.39 $\pm$ 0.02 0.66 $\pm$ 0.01 0.73 $\pm$ 0.01 0.53 $\pm$ 0.01 0.63 $\pm$ 0.01 0.67 $\pm$ 0.01 0.56 $\pm$ 0.03 0.39 $\pm$ 0.01 0.72 $\pm$ 0.02 0.64 $\pm$ 0.01 0.43 $\pm$ 0.01 0.74 $\pm$ 0.02 0.59 $\pm$ 0.01 3.42
LinOT 0.39 $\pm$ 0.02 0.64 $\pm$ 0.02 0.7 $\pm$ 0.01 0.49 $\pm$ 0.02 0.59 $\pm$ 0.0 0.62 $\pm$ 0.01 0.54 $\pm$ 0.02 0.38 $\pm$ 0.01 0.72 $\pm$ 0.01 0.63 $\pm$ 0.01 0.43 $\pm$ 0.01 0.75 $\pm$ 0.03 0.57 $\pm$ 0.01 6.54
MMD-LS 0.34 $\pm$ 0.02 0.63 $\pm$ 0.02 0.7 $\pm$ 0.01 0.46 $\pm$ 0.02 0.6 $\pm$ 0.01 0.62 $\pm$ 0.01 0.5 $\pm$ 0.02 0.36 $\pm$ 0.02 0.71 $\pm$ 0.01 0.63 $\pm$ 0.02 0.38 $\pm$ 0.01 0.74 $\pm$ 0.02 0.56 $\pm$ 0.02 6.46
Subspace JPCA 0.34 $\pm$ 0.02 0.63 $\pm$ 0.02 0.7 $\pm$ 0.01 0.46 $\pm$ 0.02 0.6 $\pm$ 0.01 0.63 $\pm$ 0.01 0.5 $\pm$ 0.02 0.35 $\pm$ 0.02 0.71 $\pm$ 0.01 0.63 $\pm$ 0.01 0.38 $\pm$ 0.01 0.74 $\pm$ 0.02 0.56 $\pm$ 0.01 10.58
SA 0.39 $\pm$ 0.02 0.64 $\pm$ 0.01 0.7 $\pm$ 0.01 0.5 $\pm$ 0.02 0.51 $\pm$ 0.21 0.63 $\pm$ 0.01 0.54 $\pm$ 0.02 0.38 $\pm$ 0.02 0.71 $\pm$ 0.02 0.63 $\pm$ 0.01 0.41 $\pm$ 0.0 0.75 $\pm$ 0.03 0.57 $\pm$ 0.03 6.50
TCA 0.02 $\pm$ 0.0 0.01 $\pm$ 0.0 0.02 $\pm$ 0.0 0.02 $\pm$ 0.01 0.02 $\pm$ 0.0 0.02 $\pm$ 0.01 0.01 $\pm$ 0.01 0.02 $\pm$ 0.01 0.02 $\pm$ 0.01 0.03 $\pm$ 0.01 0.02 $\pm$ 0.0 0.02 $\pm$ 0.01 0.02 $\pm$ 0.01 19.17
Other JDOT NA NA NA NA NA NA 0.48 $\pm$ 0.01 NA 0.65 $\pm$ 0.02 NA NA NA 0.56 $\pm$ 0.02 16.75
OTLabelProp 0.32 $\pm$ 0.02 0.57 $\pm$ 0.01 0.66 $\pm$ 0.02 0.51 $\pm$ 0.02 0.63 $\pm$ 0.01 0.63 $\pm$ 0.02 0.53 $\pm$ 0.01 0.39 $\pm$ 0.01 0.7 $\pm$ 0.01 0.63 $\pm$ 0.02 0.43 $\pm$ 0.01 0.74 $\pm$ 0.02 0.56 $\pm$ 0.01 9.25
Table 10: Accuracy score for OfficeHome dataset for each shift compared for all the deep DA methods. A white color means the method does not increase the performance compared to Train Src (Train on the source). Green indicates that the performance improved with the DA methods. The darker the color, the more significant the change.
[…]
Train Src 0.42 0.62 0.75 0.53 0.6 0.62 0.52 0.31 0.73 0.67 0.41 0.76 0.58 4.83
Train Tgt 0.78 0.91 0.86 0.8 0.93 0.85 0.8 0.78 0.86 0.78 0.76 0.92 0.83 1.00
DANN 0.44 0.63 0.73 0.58 0.61 0.62 0.54 0.38 0.75 0.67 0.44 0.76 0.6 4.25
DeepCORAL 0.47 0.64 0.75 0.63 0.59 0.65 0.59 0.39 0.76 0.7 0.45 0.78 0.62 3.00
DeepJDOT 0.47 0.65 0.74 0.63 0.65 0.66 0.59 0.44 0.77 0.71 0.47 0.79 0.63 2.33
Table 11: Accuracy score for 20NewsGroups dataset for each shift compared for all the methods. A white color means the method does not increase the performance compared to Train Src (Train on the source). Green indicates that the performance improved with the DA methods. The darker the color, the more significant the change.
[…]
Train Src 0.52 $\pm$ 0.03 0.56 $\pm$ 0.05 0.54 $\pm$ 0.05 0.7 $\pm$ 0.01 0.52 $\pm$ 0.05 0.71 $\pm$ 0.03 0.59 $\pm$ 0.04 17.58
@@
NN RW 0.52 $\pm$ 0.06 0.57 $\pm$ 0.06 0.54 $\pm$ 0.03 0.7 $\pm$ 0.01 0.52 $\pm$ 0.06 0.71 $\pm$ 0.03 0.59 $\pm$ 0.04 16.42
MMDTarS 0.52 $\pm$ 0.03 0.57 $\pm$ 0.05 0.54 $\pm$ 0.05 0.7 $\pm$ 0.01 0.52 $\pm$ 0.06 0.71 $\pm$ 0.03 0.59 $\pm$ 0.04 14.17
Mapping CORAL 0.61 $\pm$ 0.03 0.79 $\pm$ 0.02 0.62 $\pm$ 0.01 0.76 $\pm$ 0.01 0.78 $\pm$ 0.01 0.78 $\pm$ 0.02 0.73 $\pm$ 0.02 11.67
MapOT 0.67 $\pm$ 0.02 0.84 $\pm$ 0.01 0.68 $\pm$ 0.02 0.78 $\pm$ 0.02 0.83 $\pm$ 0.01 0.79 $\pm$ 0.02 0.76 $\pm$ 0.02 7.75
@@
ClassRegOT 0.92 $\pm$ 0.14 0.93 $\pm$ 0.1 0.98 $\pm$ 0.01 0.96 $\pm$ 0.01 0.99 $\pm$ 0.0 0.97 $\pm$ 0.0 0.96 $\pm$ 0.04 2.67
LinOT 0.72 $\pm$ 0.02 0.92 $\pm$ 0.0 0.72 $\pm$ 0.01 0.79 $\pm$ 0.11 0.91 $\pm$ 0.01 0.83 $\pm$ 0.07 0.82 $\pm$ 0.04 6.50
MMD-LS 0.99 $\pm$ 0.0 0.98 $\pm$ 0.01 0.99 $\pm$ 0.0 0.95 $\pm$ 0.01 0.98 $\pm$ 0.01 0.96 $\pm$ 0.01 0.98 $\pm$ 0.01 3.00
Subspace JPCA 0.71 $\pm$ 0.04 0.72 $\pm$ 0.06 0.73 $\pm$ 0.05 0.7 $\pm$ 0.01 0.84 $\pm$ 0.02 0.71 $\pm$ 0.03 0.74 $\pm$ 0.04 11.75
SA 0.83 $\pm$ 0.03 0.84 $\pm$ 0.01 0.85 $\pm$ 0.01 0.87 $\pm$ 0.01 0.84 $\pm$ 0.03 0.87 $\pm$ 0.01 0.85 $\pm$ 0.02 3.83
TCA 0.59 $\pm$ 0.25 0.52 $\pm$ 0.08 NA NA 0.55 $\pm$ 0.09 0.56 $\pm$ 0.01 0.56 $\pm$ 0.11 15.50
Other JDOT 0.67 $\pm$ 0.02 0.84 $\pm$ 0.01 0.68 $\pm$ 0.02 0.79 $\pm$ 0.02 0.84 $\pm$ 0.01 0.8 $\pm$ 0.01 0.77 $\pm$ 0.01 8.50
@@
DASVM 0.59 $\pm$ 0.16 0.77 $\pm$ 0.02 0.56 $\pm$ 0.11 0.78 $\pm$ 0.01 0.84 $\pm$ 0.05 0.79 $\pm$ 0.02 0.72 $\pm$ 0.06 10.00
Table 12: Accuracy score for AmazonReview dataset for each shift compared for all the methods. A white color means the method does not increase the performance compared to Train Src (Train on the source). Green indicates that the performance improved with the DA methods. The darker the color, the more significant the change.